Write the slope-intercept equation for the line that passes through (-2, 7) and (-7, 12).
Answer by mananth(16949)

x1 y1 x2 y2
-2 7 -7 12

slope m =(y2-y1)/(x2-x1)
(12-7)/( -7-(-2)
(5/-5)
m=-1

Plug value of the slope and point (-2,7) in
Y =mx+b
7.00=2+b
b=7-2
b=5
So the equation will be
Y =-1x+5
m.ananth@hotmail.ca